OpenStudy (anonymous):

add & simplify 2x+6/(x+3)^2 + x/x+3. a: 3x+6/(x+3)^2 b: x+2/x+3 c: x+2/(x+3)^2 d: 2/3

OpenStudy (anonymous):

@agent0smith @kelliegirl33 @Compassionate @charlotte123 please help!!!!!

OpenStudy (anonymous):

\[ \frac{(2x+6)}{(x+3)^2} + \frac{x}{(x+3)}\] So can you factor anything out of the 2x+6?

OpenStudy (anonymous):

x and 3?

OpenStudy (anonymous):

Yup yup! So then you have \[ \frac{2(x+3)}{(x+3)^2} + \frac{x}{(x+3)}\] What else can you do to that first term?

OpenStudy (anonymous):

can u cancel out the (x+3) in the numerator and denominator?

OpenStudy (anonymous):

You absolutely can ^_^ So that leaves you with \[ \frac{2}{(x+3)} + \frac{x}{(x+3)}\] Now the two terms have the same denominator, so what can you do to them?

OpenStudy (anonymous):

is the answer x+2/x+3?

OpenStudy (anonymous):

correct! ^_^

OpenStudy (anonymous):

yay! thank you!!!!!!

OpenStudy (anonymous):

Nice job ^_^ Welcome.